OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x^2 − 4x + y^2 + 8y = −4

OpenStudy (alexandervonhumboldt2):

first you need to convert this equation to (x-h)^2+(y-k)^2=r^2 where r is radius

OpenStudy (anonymous):

I know that 4 is the radius but I just cant find the rest of it.

OpenStudy (alexandervonhumboldt2):

x^2 − 4x + y^2 + 8y = −4 (x^2-4x) + (y^2-6y)=-4 (x^2-4x+4)+(y^2-6y+9)=-4+4+9 (x-2)^2+(y-3)^2=9

OpenStudy (alexandervonhumboldt2):

thus the radius is sqrt9=3, and the center is (2, 3)

OpenStudy (alexandervonhumboldt2):

(x-h)^2+(y-k)^2=r^2 where r is radius and (h, k) is center

OpenStudy (anonymous):

Oh.. Thank you for explaining it to me and working it out :).

OpenStudy (alexandervonhumboldt2):

wait i made amistake

OpenStudy (alexandervonhumboldt2):

wait

OpenStudy (anonymous):

Oh. okay!

OpenStudy (alexandervonhumboldt2):

it should be: x^2 − 4x + y^2 + 8y = −4 (x^2-4x)+(y^2+8y)=-4 (x^2-4x+4)+(y^2+8y+16)=-4+16+4 (x-2)^2+(y+4)^2=16 thus the center is (2, -4) and radius is sqrt16=4

OpenStudy (anonymous):

Thank you! :)